OpenStudy (anonymous):

what is the antiderivitive of cos(x)*sin^5(x)

zepdrix (zepdrix):

\[\large \int\limits \cos x \sin^5x \;dx\] Hmm are you familiar with the substitution method yet? It will make quick work of this problem. :) Let \(\large u=sin x\)

OpenStudy (anonymous):

you can take just sin x without the ^5?

zepdrix (zepdrix):

\[\large \sin^5x= (\sin x)^5\]Yes we want to just take the sin x as our our \(\large u\).

OpenStudy (anonymous):

OH!! Thanks! I got it now